(9.3.2) Listed below are body temperatures from five different subjects measured at 8 AM and again at 12 AM. Find the values of d and sd. In​ general, what does alpha d ​represent?
Temperature (F) at 8am 98.4,98.8,97.3.,97.3,97.3,97.6
Temperature (F) AT 12am 98.8,99.1,97.7,97.0,98.1
d=?
1
Expert's answer
2020-11-17T14:05:48-0500

From above given data, following gives the difference values :


Diff (T2 - T1)


0.4

0.3

0.4

-0.3

0.5


Formula to find sd:


"\sum{x} = 1.3" "{\sum}{x^2} = 0.75"


"{\sum}{x^2} - \frac{({\sum}{x})^2}{n} = 0.75 - 0.338 = 0.412"


"{s^2}= \frac{{\sum}{x^2} - \frac{({\sum}{x})^2}{n}}{n-1}
 = \frac{ 0.412 }{4} = 0.103"


"{s} = \sqrt{0.103} = 0.3209"


Alpha represents significance level